OpenStudy (anonymous):

a hungry wolf picks up a 15.95N rabbit in its mouth and lifts it 0.575m. it then carries the rabbit 15.0m to its den. what is the work being done? last question i need for my module, pls

OpenStudy (anonymous):

W=F*s W1=15,95N*0,575m=9,17 J W2=15,95 N * 15m =239,25J W=W1+W2=248,42 J